I am having problems with finding the center and radius.
Here is the problem : Find the center and radius of the circle given in general form x^2+y^2-10x-2y+50=49
Could someone help me with this?

Answer by stanbon(75887)

Find the center and radius of the circle given in general form
x^2+y^2-10x-2y+50=49
Complete the square on the x terms and on the y terms:
x^2-10x + ? + y^2 - 2y + ? = -1
---
x^2 - 10x + 5^2 + y^2 - 2y + 1^2 = -1 + 5^2 + 1^2
---
Factor:
(x-5)^2 + (y-1)^2 = 25
============================
Center at (5,1) ; Radius = sqrt(25) = 5
================================================
Cheers,
Stan H.